OpenStudy (shamil98):

yo my real niqq3r @Luigi0210 and my niqer @awkwardpanda

OpenStudy (awkwardpanda):

XD "niqer"

OpenStudy (shamil98):

okie bud lets start .

OpenStudy (shamil98):

we left off at factoring yesterday for finding limits.

OpenStudy (awkwardpanda):

yus :3

OpenStudy (shamil98):

let's see if you held the info. evaluate the limit \[\lim_{x \rightarrow 2} \frac{ 2x^2 - 3 }{ x + 5 }\]

OpenStudy (awkwardpanda):

\[\frac{ 2(2)^2-3}{ 2+5 }\]

OpenStudy (shamil98):

yer

OpenStudy (shamil98):

simplify

OpenStudy (awkwardpanda):

okai

OpenStudy (awkwardpanda):

7? >.>

OpenStudy (shamil98):

\[\frac{ 5 }{ 7 }\]...

OpenStudy (awkwardpanda):

DENG IT it's the simple things that i forget ;-;

OpenStudy (shamil98):

ok, you can do that later, first lets learn about discontinuities and then the conjugation method

OpenStudy (awkwardpanda):

okay :3

OpenStudy (shamil98):

do you know how to find the domain of a rational function? or a function in general? this is a rational function \[f(x) = \frac{ x^2 - 5 }{ x+5 }\]

OpenStudy (awkwardpanda):

could you remind me pls .-.

OpenStudy (shamil98):

ok, so you don't know.. let's start with the rational function.. you set the denominator to equal zero and solve .. \[x + 5 = 0\] x = -5 the domain is all real numbers except -5, because the denominator becomes 0 and you can't divide by zero. \[[D: \mathbb{R} : x \ne -5]\]

OpenStudy (shamil98):

well that's not the correct notation for writing domain but w/e ill do that later.

OpenStudy (shamil98):

its something like that i forget.

OpenStudy (awkwardpanda):

i get it, in a way XD

OpenStudy (shamil98):

anyways -5 is the domain , the vertical asymptote where this function does not touch  the graph never touches -5